OpenStudy (anonymous):

Second degree Taylor Series of ln(x) at x=2?

OpenStudy (anonymous):

I ended up with \[\ln 2 + \frac{ 1 }{ 2 } (x-2) - \frac{ 1 }{ 4 } (x-2)^{2}\]

OpenStudy (anonymous):

Which I don't believe is right. Supposedly, the 1/4 should be 1/8, and I don't know how.